OpenStudy (anonymous):

The table below shows the fees to check bags on a certain airline. A 2-column table with 'Type of Bag' and 'Fee per Bag' as the column headings. The first checked bag costs $20, the second checked bag costs $25, and extra bags cost $50 each If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (anonymous):

20 + 25 + 50(x - 2) = C = 145 Expand: 20+25+50x-100 =145 50x-55=145 50x = 200 x =?

OpenStudy (anonymous):

4